in Exhibits t0s E10.24 Straightforward calculation of variances; variance diagrams: manufacturer o(12 ca Apollo Can Ltd manufactures recyclable soft drink cans. A unit of production is a box he tollowing standards have been set by the production engineering staff and the factory accountant Direct materia Quantity: 4 kilograms Price:$0.60 per kilogram Direct labour 0.62-. Quantity: 0.25 hour Rate: $20 per hour 0bx 2opo- Actual material urchases amounted to 240000 kilograms af $0.62 per kilogram Actual costs incurred in the production of 50000 units were as follows: Direct material $130200 for 210000 kilograms $273000 for 13000 hours Direct labour Required: 1 Construct an Exelw spreadsheet to calculate the direct material price and quantity varía and the direct labour rate and efficiency variances.Indicate whether each variance is fa or unfavourable. Use your spreadsheet to show how the solution will change if the following informatio standard direct material price is $0.59 per kilogram and standard direct labour rate is vou rab 2
Answer #1

1)Direct Material Price Variance = (Standard Price- Actual Price)x Actual Quantity

=(0.60- 0.62) x 210000

=4,200 (Unfavorable)

Direct Material Quantity Variance = (Standard Quantity- Actual Quantity) x Standard Price

=(50,000x 4 - 210000)0.60

=6,000 (Unfavorable)

Direct Labor Rate Variance = (Standard Rate- Actual rate)x Actual Hours

=(20 -21) x 13000

13,000 (Unfavorable)

Direct Labor Quantity Variance = (Standard Hours - Actual Hours ) x Standard Rate

= (50,000 x 0.25 -13,000)x 20

=10,000 (Unfavorable)

2)

Direct Material Price Variance = (Standard Price- Actual Price)x Actual Quantity

=(0.59- 0.62) x 210000

=6,300 (Unfavorable)

Direct Material Quantity Variance = (Standard Quantity- Actual Quantity) x Standard Price

=(50,000x 4 - 210000)0.59

=5,900 (Unfavorable)

Direct Labor Rate Variance = (Standard Rate- Actual rate)x Actual Hours

=(22 -21) x 13000

13,000 (favorable)

Direct Labor Quantity Variance = (Standard Hours - Actual Hours ) x Standard Rate

= (50,000 x 0.25 -13,000)x 22

=11,000 (Unfavorable)